OpenStudy (anonymous):

Basic multiplication !!! ln(1+r^2)rdrd(theta) So i have to multiply ln(1+r^2) by r dont want to make silly mistakes....

OpenStudy (anonymous):

Are you sure this is multiplications? This looks more like integration from calculus with multivariables

OpenStudy (anonymous):

It a double integral question with polar coordinates\[\int\limits_{0}^{5}\int\limits_{0}^{\sqrt{25-x^{2}}}\ln \left( 1+x^{2}+y^{2} \right)dydx\]

OpenStudy (anonymous):

and i ended up with what i posted when trying to change to polar coordinates.... does that make sense ? hope im not doing it wrong :/

OpenStudy (anonymous):

So in your second, post that is the original equation? and when you switched it to polar coordinates you got the equation in your opening post?

OpenStudy (anonymous):

Yes thats correct , with the limits being \[\int\limits_{0}^{\frac{ \pi }{ 2 }}\int\limits_{0}^{5}\]

OpenStudy (anonymous):

First one is pi/2 if its too small to see :)

OpenStudy (anonymous):

So do i just take r to the front \[rln \left( 1+r^2{?} \right)\]

OpenStudy (anonymous):

and then intergrate ?

OpenStudy (anonymous):

Give me a second, I haven't done these problems in the longest time. lol. If someone else comes by before me they can give it a shot, trying to relearn this real quick

OpenStudy (anonymous):

No worries :) i appreciate your help :)

OpenStudy (anonymous):

Have you ever heard of "u" substitution? Let u = 1 + r^2

OpenStudy (anonymous):

Yes i have

OpenStudy (anonymous):

Do you have the answer from teh back of the book or something i wanna check something

OpenStudy (anonymous):

No i dont sorry :(

OpenStudy (anonymous):

okay, so you know to work from the inside out from your polar coordinate double integral equation, do the "u" sub I told you, since we can't multiply "r" into ln(1+r^2).

OpenStudy (anonymous):

so we get.... rln(u) where u =1+r^2 du/dr=2r dr=du/2r ????

OpenStudy (anonymous):

Yeah. The odd thing is, I'm getting a negative answer which is incorrect because integration over double integrals usually provide and area with positive value not negative. interesting. are you sure the limits are correct?

OpenStudy (anonymous):

hmmm.. they could be wrong ill show you what i did

OpenStudy (anonymous):

\[y^{2}\le \sqrt{25-x^{2}}\]\[x^{2}+y^{2}\le25\]\[x \ge0 , x \le5\]

OpenStudy (anonymous):

okay, that works for the limits of "r" what about "theta"

OpenStudy (anonymous):

I havent got that far yet... what answer did u get once you intergrated it??

OpenStudy (anonymous):

Final Answer as\[\frac{ -25 }{ 52 }\frac{ \pi }{ 2 }\]

OpenStudy (anonymous):

Sorry i havent got that far yet...im still trying to intergrate with sub !! its been a while since i have done it

OpenStudy (anonymous):

Yeah "u" subs will come back and haunt you. but let me ask someone real quick before you start going with the "u" subs

OpenStudy (anonymous):

@ganeshie8 Can you help us a bit here ganeshie8

OpenStudy (anonymous):

I agree!! parts all the way!!

OpenStudy (anonymous):

@wio Need your expertise real quick, second pair of eyes please.

OpenStudy (anonymous):

I typed the function into wolf frame alpha and got a circle that had its volume under the xy plane, maybe that's why it's negative, but we'll see from the other experts here.

OpenStudy (anonymous):

Hmm so does that mean my limits are wrong ?

OpenStudy (anonymous):

No. Sometimes the answers are negative. Is this from your teacher or a textbook?

OpenStudy (anonymous):

https://www.google.com/#q=graph+ln%281%2Bx^%282%29%2By^%282%29%29 that's what it looks like.

OpenStudy (anonymous):

its a question of a previous years exam, but they haven't provided solutions to the problem

OpenStudy (anonymous):

Oh. Interesting. Well, it's best to wait for another pair of eyes to see this before you go through, if else, skip this one and study another thing while you wait for confirmation on this. @Directrix Sorry I couldn't fully help you.

OpenStudy (anonymous):

Yes plenty of other things to do.... dont suppose your interested in a eigenvalue and eigenvector problem??

OpenStudy (anonymous):

Thanks for your help !! it got me thinking more having someone to talk to !!

OpenStudy (anonymous):

You're welcome. I'm going to sleep now it's late over here. Good night.

OpenStudy (anonymous):

Okay then. Thank you once again!!!